OpenStudy (samsan9):

rewrite the following expression as a single logarithm 2log(x+3)+3log(x-7)-5log(x-2)-log(x^2)

OpenStudy (wearehiding):

Use this first: \(\Large alog(x+b)=log(x+b)^a\)

OpenStudy (samsan9):

its multiple choice if that helps i could try to take a pic of the answers

OpenStudy (mathmate):

and \(\large log(x+a)+log(x+b)=log((x+a)(x+b))\)

OpenStudy (samsan9):

so i know that is going to be in fraction form

OpenStudy (mathmate):

Well, with multiple choice problems, it's best not to look at the answer, work out the answer, and then compare. This will test your knowledge on the subject and give you confidence during exams. Multiple choice questions are more difficult to score than regular problems because most people spend less time and think they can get the answer without writing anything down.